OpenStudy (anonymous):

Which of the following rectangular prisms has a base area of 40 cm2 and a volume of 160 cm3?

OpenStudy (anonymous):

Select one: a. Height = 10 cm, base length = 4 cm, base width = 4 cm b. Height = 2 cm, base length = 10 cm, base width = 4 cm c. Height = 4 cm, base length = 10 cm, base width = 4 cm d. Height = 4 cm, base length = 4 cm, base width = 4 cm

OpenStudy (calculusxy):

Base area would be \(\large length \times width\) .

OpenStudy (calculusxy):

Volume is calculated by \(\large \text{Length} \times \text{Width} \times \text{Height}\)

OpenStudy (calculusxy):

Which one do you think will be the answer, given the formulas above?

OpenStudy (anonymous):

Length x Width

OpenStudy (calculusxy):

That's not an answer choice, I hope you know.

OpenStudy (anonymous):

Yeah i think its A though

OpenStudy (calculusxy):

Check this: length = 4 width = 4 is \(4 \times 4\) equal to the base area of \(40 cm^2\)

OpenStudy (anonymous):

I thought it was 4x10= 40 And then 4x4=16 , 16x10=160

OpenStudy (calculusxy):

Not really, you must remember to understand what each measurement really represents. So you must look closely in this case.

OpenStudy (calculusxy):

Height = 10 cm, base length = 4 cm, base width = 4 cm Notice that only the \(\large \text{Height}\) is \(\large 10 cm\). The \(\large \text{Base Length}\) is \(\large 4 cm\). The \(\large \text{Base Length}\) is \(\large 4 cm\). So you MUST watch out!

OpenStudy (anonymous):

So then its C

OpenStudy (calculusxy):

Correct!

OpenStudy (anonymous):

Thanks for the help

OpenStudy (calculusxy):

You're welcome :)